Differentiate each function given below with respect to x.

(i) f(x) = √tan^−1x

(ii) f(x) = cos^−1(e^2x)

(iii) f(x) = sin^−1(√1− √x^2 )

(iv)f(x) = tan^−1(x/√x-x^2)

(v) f(x) = sin( tan^−1 4x)
1
Expert's answer
2020-06-01T19:46:44-0400

(i) "f(x)=\sqrt{\tan^{-1}(x)}"

Differentiating with respect to x both sides,

"f'(x) = [\frac{1}{2\sqrt {tan^{-1} x}}] [\frac{1}{1 + x^2}]"


(ii) f(x) = cos-1(e2x)

Differentiating with respect to x both sides,

"f'(x) = [\frac{-2e^{2x}}{\sqrt{1 - {e^{4x} }}} ]"


"(iv) f(x) = tan^{-1}\frac{x}{\sqrt{x - x^{2}}}"

Differentiating with respect to x both sides,

"f'(x) = [\frac{1}{1 + {\frac{x^{2}}{x - x^{2}} }} ] [\frac{\sqrt{x - x^{2}} - \frac{x{(1 - 2x)}}{2\sqrt{x - x^{2}}}}{x - x^{2}}]"


Solving equation, we get,

"f'(x) = \frac{(1-x)(2x - 2x^{2} - x + 2x^{2})}{2 {(x - x^{2})}^{3/2}}"


"f'(x) = \frac{1}{2 } [ ({x - x^{2}})^{(- {\frac{1}{2}})}]"


"(iii) f(x) = {sin^{-1}{\sqrt{1 - \sqrt{x^{2} }}}}"

"so, f(x) will be \implies f(x) = {sin^{-1}{\sqrt{1 - x}}}"

Differentiating both sides w.r.t. x


"f'(x) = \frac{1}{\sqrt{1 - (1-x)}} \frac{-1}{(2\sqrt{1-x})}"

"f'(x) = \frac{-1}{2 \sqrt{x- x^{2}}}"


(v) F(x) = sin(tan-14x)

Differentiating both sides with respect to x,

"f'(x) = \frac{(4 cos(tan^{-1}4x))}{1 + 16x^{2}}"